# Vacancy-Induced Quantum Properties in 2D Silicon Carbide: Atomistic insights from semi-local and hybrid DFT calculations

### v1

@@ I Introduction
Moreover, DFT calculations have provided valuable insights into the spin-strain coupling properties of defects in 2D SiC [28]. By analyzing the electronic structure and spin density distribution of defects under different strain conditions, researchers can predict the sensitivity of defect-related optical transitions to mechanical deformation. This knowledge is crucial for developing strain-tunable quantum emitters and sensors based on 2D SiC. Recent theoretical studies have also explored the potential of defect engineering in 2D SiC through the incorporation of foreign atoms or the creation of specific defect complexes [29]. DFT calculations have predicted that doping 2D SiC with elements such as nitrogen or boron can introduce new energy levels and optical transitions, enabling the tuning of the material’s electronic and optical properties. Recent experimental breakthroughs in synthesizing 2D SiC have validated many theoretical predictions about its properties and stability. Both top-down chemical exfoliation [30] and bottom-up epitaxial growth on transition metal carbides [31] have demonstrated the feasibility of producing monolayer SiC. These advances provide crucial experimental validation of the structural and electronic properties predicted by theory, while also revealing new phenomena such as strong spin-orbit coupling effects in substrate-supported 2D SiC [31]. However, the accuracy of DFT calculations hinges on the chosen exchange-correlation functional. While PBE offers an efficient description of electron localization, it often falls short in capturing the correct electronic structure and defect formation energies [32, 33, 34]. PBE tends to underestimate the band gap of semiconductors and overdelocalize the defect states, leading to an underestimation of defect formation energies and charge transition levels. This limitation can be particularly problematic for defects in wide-bandgap materials like SiC, where the accurate description of localized defect states is crucial. Alternatively, hybrid functionals like HSE06 have emerged as a more reliable choice for defect calculations in semiconductors [35, 36, 37]. By incorporating a fraction of exact exchange from Hartree-Fock theory, HSE06 can partially correct the self-interaction error and improve the description of localized defect states. HSE06 has been shown to provide accurate band gaps and defect formation energies for a wide range of semiconductor materials, including 2D SiC. However, the computational cost of HSE06 calculations is significantly higher than that of PBE, which can limit its applicability to large-scale simulations. To bridge the gap between accuracy and computational efficiency, meta-GGA functionals like SCAN and r2SCAN have been developed [38, 39, 40, 41]. These functionals go beyond the local density approximation (LDA) and GGA by incorporating higher-order derivatives of the electron density, such as the kinetic energy density. SCAN and r2SCAN have shown promising performance in describing the electronic structure and thermodynamic properties of a wide range of materials, including semiconductors and 2D systems. Recent studies have demonstrated that SCAN and r2SCAN can provide accurate band gaps and defect formation energies for 2D SiC, comparable to those obtained with HSE06 [40, 41]. Moreover, these functionals have been shown to correctly capture the localization of defect states and the charge transition levels, which are crucial for understanding the optical and transport properties of defects in 2D SiC. The computational cost of SCAN and r2SCAN calculations is significantly lower than that of HSE06, making them attractive options for large-scale simulations of defects in 2D materials. However, it is important to note that while meta-GGA functionals like SCAN and r2SCAN offer improved accuracy compared to PBE, they are not a universal solution for all materials and properties. In some cases, the inclusion of exact exchange through hybrid functionals like HSE06 may still be necessary to correctly describe the electronic structure and defect properties of 2D SiC. Therefore, a careful benchmarking of different functionals against experimental data and high-level theoretical methods is essential to ensure the reliability of DFT calculations for defects in 2D SiC. The objective of this research is to examine the formation energies of Si and C vacancies in bulk and monolayer SiC using various DFT functionals. The primary aim is to evaluate the effectiveness of each functional in representing localized defect states and to comprehend the impact of these states on the material’s quantum characteristics. The study aims to provide valuable insights into the theoretical basis of SiC’s defects and aid in selecting appropriate computational strategies in the rapidly expanding field of SiC-based technologies. We will delve into the electronic structures of these defects and establish a correlation between their fundamental electronic properties and emerging quantum features. The ensuing sections will present the findings of this study, covering the defect formation energies (Fig. 1), electronic structure (Fig. 3), vacancy migration (Fig. 5), and optical properties (Fig. 6) of silicon and carbon vacancies in monolayer SiC.
II Computational Methods
We conducted spin-polarized DFT calculations using the Vienna Ab initio Simulation Package (VASP) [42] with plane-wave basis sets, PAW-PBE pseudopotentials [43], and an energy cutoff of 450 eV. We employed the Generalized Gradient Approximation (GGA) with Perdew-Burke-Ernzerhof (PBE) [44], the strongly constrained and appropriately normed (SCAN) meta-GGA type [45] and its modified version r2SCAN [46], as well as the Heyd-Scuseria-Ernzerhof (HSE06) hybrid exchange-correlation functionals [47], including DFT-D3 van der Waals correction [48]. Complete relaxation of atomic positions and supercell dimensions was allowed using conjugate gradient optimization techniques without symmetry constraints. Force convergence criteria of 0.02 eV/ and total energy precision within 0.01 meV were achieved with Brillouin zone sampling on a 221 Monkhorst-Pack k-point grid for structural optimizations and a denser 441 grid for band structure calculations [49]. Migration pathways and diffusion barriers () of defect sites were computed using the climbing-image nudged elastic band (CI-NEB) method [50], as presented in Fig. 5. The optical properties were calculated using density functional perturbation theory (DFPT) within the PBE framework, which provides a computationally efficient approach for excited state properties while maintaining good accuracy for relative changes in optical response. The absorption coefficients () were computed from the complex dielectric function and are presented as percentages relative to the pristine case to facilitate direct comparison of vacancy effects. The energy range was chosen to focus on the far-infrared region (0-0.5 eV) where vacancy-induced features are most prominent. We determined vacancy defect formation energies in monolayer SiC by applying thermodynamic principles, considering carbon-rich environment. In this work, is derived from diamond silicon, while is calculated from bulk graphite. These calculations provide insight into the preference for different vacancy defects under varied experimental conditions, as shown in Fig. 1. The Freysoldt, Neugebauer, and Van de Walle (FNV) [51] and Kumagai & Oba (KO) [52] methods are used to correct for electrostatic interactions in the calculation of defect formation energies in charged systems. The formation energies of defects in charge-state were calculated using the equation:
Here, is the total energy of the supercell containing one defect of charge state , is the total energy of the defect-free supercell, is the chemical potential of the removed atom type, is the number of missing atoms, is the Fermi energy referenced to the theoretical valence band maximum (VBM) for a given functional, and is the correction term for electrostatic interactions between periodically repeated defects, calculated using the FNV or KO method. Freysoldt’s method, which is often employed in defect calculations, addresses the spurious electrostatic interaction between periodic images of charged defects in a supercell. In this scheme, a model potential is used as the basis for canceling the long-range Coulomb interaction as well as aligning the potential with the potential of a perfect bulk crystal numerically.
where is the charge of the defect, is the dielectric constant, and is the supercell dimension. The correction term consists of two parts: the alignment of the potential, , and the image charge correction, which is a function of the charge , the dielectric constant , and the supercell size . The potential alignment term ensures that the defect level aligns with the bulk level far from the defect. Physically, this method acknowledges that the potential due to a charged defect should not differ from the perfect crystal at a sufficiently large distance. This aligns the electrostatic potential outside the defect region with that of the bulk, ensuring that the calculated defect formation energy reflects the intrinsic energy without artificial stabilization from periodic images. On the other hand, the Kumagai method extends the Freysoldt approach by including the anisotropic nature of the dielectric response. It uses the full dielectric tensor and computes the potential alignment term using anisotropic dielectric models, which is crucial for materials with directional bonding, such as SiC.
The anisotropic dielectric constant replaces the isotropic used in Freysoldt’s method. The corrected formation energy takes into account the varying dielectric response depending on the direction relative to the crystal axes. Here, and are computed considering the full tensor, which reflects the anisotropic electronic screening response of the material. The optical properties were calculated using density functional perturbation theory (DFPT) within the PBE framework, which provides a computationally efficient approach for excited state properties while maintaining good accuracy for relative changes in optical response. The absorption coefficients () were computed from the complex dielectric function () and are presented as percentages relative to the pristine case to facilitate direct comparison of vacancy effects. The energy range was chosen to focus on the far-infrared region (0-0.5 eV) where vacancy-induced features are most prominent.
III Results and Discussion
III.1 Defect Formation Energies
[…]

### v2

@@ 1 Introduction
The field of two-dimensional (2D) semiconductors has witnessed remarkable progress in recent years, driven by the discovery of novel materials and their intriguing properties 1, 2, 3. Following the initial isolation of graphene, other 2D materials, including transition metal dichalcogenides (TMDs), hexagonal boron nitride (h-BN), and phosphorene, have been extensively investigated. These atomically thin materials, characterized by their unique electronic, optical, and mechanical properties, have opened exciting avenues for applications, particularly in quantum information science and optoelectronics 4, 5, 6, 7, 8, 9, 10. Among these emerging 2D semiconductors, silicon carbide (SiC) has attracted significant attention due to its exceptional quantum properties that persist even at room temperature 7, 1, 2. Defects within SiC, such as silicon vacancies (V) and carbon antisites (Si), play a pivotal role in enabling single-photon emission, a critical feature for quantum communications 1, 2, 3. Through advanced spectroscopic techniques like electron spin resonance (ESR) and photoluminescence (PL), researchers have gained valuable insights into the nature of these defect states and their impact on the material’s electronic and optical properties 4, 5, 6, 7, 8, 9, 10. Theoretical predictions suggest that local defects in 2D SiC could offer superior photon emission efficiency due to reduced dielectric screening and enhanced quantum confinement effects 11, 12. The single-atomic-layer thickness of 2D SiC minimizes the path length for emitted photons, potentially increasing extraction efficiency compared to bulk materials where total internal reflection can trap emitted photons. Additionally, the reduced dimensionality enhances the oscillator strength of optical transitions, potentially leading to brighter single-photon sources. Furthermore, the existence of Stone-Wales defects 13, 14, created by rotating a carbon-carbon bond in the SiC lattice, has been proposed as offering superior radiative transitions compared to isolated point defects due to their unique electronic structure. Despite these promising theoretical predictions, realizing the potential of 2D SiC in practice remains a significant challenge due to difficulties in synthesizing high-quality 2D SiC and controlling defects at the atomic level.15, 16, 17, 18, 19 Current synthesis approaches include both top-down methods such as selective etching and mechanical exfoliation, and bottom-up techniques like chemical vapor deposition and molecular beam epitaxy.20, 21 However, these methods often result in samples with high defect concentrations or limited size, hindering their practical applications. Controlling defect type, concentration, and spatial distribution remains a critical challenge that must be addressed to harness the quantum properties of 2D SiC effectively. To characterize and study the defects in 2D SiC, advanced techniques with atomic-scale resolution are essential. Scanning tunneling microscopy (STM),21 high-resolution transmission electron microscopy (HRTEM), and specialized optical characterization methods like photoluminescence (PL) spectroscopy have been instrumental in probing the atomic structure and properties of defects.7, 22, 23, 24, 25, 26 Substantial progress has been achieved in comprehending the dynamics of defects in 2D SiC via advanced theoretical research, particularly with respect to thermodynamic stability.14, 12, 27 First-principles calculations based on density functional theory (DFT) have provided valuable insights into the formation energies, electronic structure, and optical properties of various defects in 2D SiC 12, 14, 27, 28. By comparing the calculated defect formation energies and charge transition levels with experimental data, researchers can identify the most stable defect configurations and predict their behavior under different environmental conditions. Pizzochero and Berthod14 conducted a comprehensive DFT study examining the energetic stability of various defects in monolayer SiC, finding that the divacancy complex (V-V) exhibits remarkably low formation energy and is thermodynamically favorable compared to isolated vacancies. This finding has important implications for experimentalists aiming to engineer and control defects in freestanding 2D SiC, as the divacancy-complex can serve as a stable and optically active center for quantum applications. Moreover, DFT calculations have provided valuable insights into the spin-strain coupling properties of defects in 2D SiC.12 By analyzing the electronic structure and spin density distribution of defects under different strain conditions, researchers can predict the sensitivity of defect-related optical transitions to mechanical deformation. This knowledge is crucial for developing strain-tunable quantum emitters and sensors based on 2D SiC. Udvarhelyi et al.12 demonstrated through DFT calculations that certain defect complexes in 2D SiC exhibit strong spin-strain coupling, potentially enabling mechanical control of their quantum states. Recent theoretical studies have also explored the potential of defect engineering in 2D SiC through the incorporation of foreign atoms or the creation of specific defect complexes.27 DFT calculations have predicted that doping 2D SiC with elements such as nitrogen or boron can introduce new energy levels and optical transitions, enabling the tuning of the material’s electronic and optical properties. Recent experimental breakthroughs in synthesizing 2D SiC have validated many theoretical predictions about its properties and stability.19, 29 Both top-down chemical exfoliation19 and bottom-up epitaxial growth on transition metal carbides29 have demonstrated the feasibility of producing monolayer SiC. These advances provide crucial experimental validation of the structural and electronic properties predicted by theory, while also revealing new phenomena such as strong spin-orbit coupling effects in substrate-supported 2D SiC.29 However, the accuracy of DFT calculations hinges critically on the chosen exchange-correlation functional. The Perdew-Burke-Ernzerhof (PBE) functional offers an efficient approximation but often falls short in capturing the correct electronic structure and defect formation energies.30, 31, 32 PBE tends to underestimate the band gap of semiconductors and overdelocalize defect states, leading to an underestimation of defect formation energies and charge transition levels. This limitation can be particularly problematic for defects in wide-bandgap materials like SiC, where accurate description of localized defect states is crucial. The self-interaction error inherent in semi-local functionals like PBE arises from the incomplete cancellation of the electron self-interaction in the exchange-correlation energy. This error is particularly pronounced for localized states, such as those introduced by defects in semiconductors, leading to an artificial delocalization of these states and consequently incorrect predictions of their energetics and electronic structure. The underestimation of bandgaps in semiconductors by PBE is a direct consequence of this self-interaction error, which affects both the valence and conduction band edges but typically lowers the conduction band more significantly, resulting in reduced bandgaps compared to experimental values. Hybrid functionals like HSE06 have emerged as a more reliable choice for defect calculations in semiconductors.33, 34, 35 By incorporating a fraction of exact exchange from Hartree-Fock theory 36, 37, HSE06 can partially correct the self-interaction error and improve the description of localized defect states. HSE06 has been shown to provide accurate band gaps and defect formation energies for a wide range of semiconductor materials, including 2D SiC. The mixing of exact exchange with semi-local exchange in HSE06 is particularly effective at correcting the self-interaction error for localized states, resulting in more accurate predictions of defect formation energies, charge transition levels, and magnetic properties. However, the computational cost of HSE06 calculations, scaling as due to the inclusion of exact exchange, is significantly higher than that of PBE, which scales as , limiting HSE06’s applicability to large-scale simulations. To bridge the gap between accuracy and computational efficiency, meta-GGA functionals like SCAN and r2SCAN have been developed.38, 39, 40, 41 These functionals go beyond the local density approximation (LDA) and GGA by incorporating higher-order derivatives of the electron density, such as the kinetic energy density. SCAN and r2SCAN have shown promising performance in describing the electronic structure and thermodynamic properties of a wide range of materials, including semiconductors and 2D systems. Recent studies have demonstrated that SCAN and r2SCAN can provide accurate band gaps and defect formation energies for 2D SiC, comparable to those obtained with HSE06.40, 41 The strongly constrained and appropriately normed (SCAN) meta-GGA functional was designed to satisfy all known exact constraints for exchange-correlation functionals and is appropriately normed to accurate reference data for systems where the exact functional should recover the exact energy.38 By incorporating the kinetic energy density, SCAN can distinguish between different chemical bonding environments, leading to improved descriptions of both covalent bonds and weak interactions. This is particularly beneficial for materials like SiC, which feature a mix of covalent and ionic bonding characters. However, SCAN has been found to be numerically challenging to converge in some cases, leading to the development of the regularized SCAN (r2SCAN) variant, which maintains most of SCAN’s accuracy while offering improved numerical stability.40 The computational cost of SCAN and r2SCAN calculations is significantly lower than that of HSE06, making them attractive options for large-scale simulations of defects in 2D materials. While these functionals offer improved accuracy compared to PBE, they may still not fully capture the exact exchange effects necessary for highly localized defect states, particularly those with strong magnetic character. Therefore, a careful comparison of different functionals is essential for establishing reliable computational protocols for studying defects in 2D SiC. The objective of this research is to systematically examine the formation energies, electronic properties, migration barriers, and optical responses of Si and C vacancies in monolayer SiC using multiple DFT functionals (PBE, SCAN, r2SCAN, and HSE06). By comparing these functionals, we aim to evaluate their effectiveness in representing localized defect states and to comprehend the impact of these states on the material’s quantum characteristics. Our comprehensive approach includes accurate charge correction schemes for charged defects and detailed analysis of spin polarization effects. The study aims to provide valuable insights into the theoretical basis of SiC’s defects and aid in selecting appropriate computational strategies in the rapidly expanding field of SiC-based technologies.
2 Computational Methods
We conducted spin-polarized density functional theory (DFT) calculations using the Vienna Ab initio Simulation Package (VASP)42 with plane-wave basis sets, PAW-PBE pseudopotentials,43 and an energy cutoff of 450 eV. This energy cutoff was chosen after convergence testing, which showed that increasing the cutoff beyond 450 eV changed total energies by less than 1 meV per atom. (see Supplementary Figure S1 for convergence graphs) We employed four different exchange-correlation functionals to systematically evaluate their performance in describing defect properties: (1) the Generalized Gradient Approximation (GGA) with Perdew-Burke-Ernzerhof (PBE),30 (2) the strongly constrained and appropriately normed (SCAN) meta-GGA functional,38 (3) its modified version r2SCAN,40 and (4) the Heyd-Scuseria-Ernzerhof (HSE06) hybrid exchange-correlation functional,33 all including DFT-D3 van der Waals corrections.44 Complete relaxation of atomic positions and supercell dimensions was allowed using conjugate gradient optimization techniques without symmetry constraints. Force convergence criteria of 0.02 eV/ and total energy precision within 0.01 meV were achieved with Brillouin zone sampling on a 2×2×1 Monkhorst-Pack k-point grid for structural optimizations and a denser 4×4×1 grid for band structure calculations.45 This k-point sampling was verified to be sufficient through convergence tests that showed energy differences less than 5 meV when compared to denser grids (Supplementary Figure S2). We employed a 3×3 supercell (containing 18 atoms for pristine SiC) for simulating the defective systems. This supercell size was chosen as a balance between computational cost and accuracy, and we performed convergence tests with larger 4×4 and 5×5 supercells for selected configurations to verify that our 3×3 supercell results are well-converged. Energy differences between the 3×3 and 5×5 supercells were found to be less than 0.05 eV for formation energies (Supplementary Figure S3) and less than 0.1 eV for migration barriers, confirming the adequacy of our chosen supercell size. The initial crystal structure of monolayer SiC was obtained from theoretical models of the hexagonal 2D SiC lattice, as reported in previous studies 28, 46. A 3×3 supercell containing 18 atoms (9 Si and 9 C) was constructed from this pristine structure. Vacancies were introduced by removing a single Si or C atom from the supercell, followed by full structural relaxation to account for local distortions around the defect site. A vacuum length of 20 in the z-direction was used to prevent spurious interactions between periodic images. This vacuum size was determined through convergence testing, which showed negligible energy changes ( 1 meV) when increasing the vacuum beyond 20 . For the electronic structure calculations, we employed the tetrahedron method with Blöchl corrections for Brillouin zone integration to obtain accurate density of states (DOS) and band structures. The DOS calculations utilized a denser 6×6×1 k-point grid to ensure smooth spectra, while band structures were calculated along the high-symmetry path in the Brillouin zone. For charged defect calculations, we included a compensating homogeneous background charge to maintain overall charge neutrality in the periodic system, as is standard practice in plane-wave DFT calculations.
We determined vacancy defect formation energies in monolayer SiC by applying thermodynamic principles and considering carbon-rich conditions. The formation energies of defects in charge-state were calculated using the equation:
Here, is the total energy of the supercell containing one defect of charge state , is the total energy of the defect-free supercell, is the chemical potential of the removed atom type, is the number of missing atoms, is the Fermi energy referenced to the theoretical valence band maximum (VBM) for a given functional, and is the correction term for electrostatic interactions between periodically repeated defects. The chemical potentials of the constituent elements are subject to thermodynamic constraints. In this work, is derived from diamond silicon, while is calculated from bulk graphite. This approach corresponds to carbon-rich conditions, which are likely to be encountered in experimental synthesis of 2D SiC. Specifically, we calculated the bulk energies of diamond silicon and graphite, and used these values as the respective chemical potentials, ensuring that (bulk) to maintain thermodynamic equilibrium with bulk SiC. These chemical potential references were chosen to represent realistic experimental conditions, though the absolute formation energies would differ under silicon-rich conditions. For charged defect calculations, we employed two different charge correction schemes to account for the spurious electrostatic interactions that arise in periodic supercell approaches: the Freysoldt, Neugebauer, and Van de Walle (FNV) method47 and the Kumagai & Oba (KO) method.48 These corrections are essential for obtaining accurate formation energies of charged defects, as the periodic boundary conditions used in plane-wave DFT calculations can lead to significant errors due to the long-range nature of Coulomb interactions.
The FNV method addresses the spurious electrostatic interaction between periodic images of charged defects in a supercell using an isotropic dielectric constant:
where is the charge of the defect, is the dielectric constant, is the supercell dimension, is the image charge correction term, and is the potential alignment term. The KO method extends the Freysoldt approach by including the anisotropic nature of the dielectric response, which is particularly important for 2D materials:
where is the anisotropic dielectric tensor. This approach accounts for the different screening responses in the in-plane and out-of-plane directions, providing more accurate corrections for charged defects in 2D systems. To obtain the dielectric properties needed for these corrections, we calculated the static dielectric constant and Born effective charges using density functional perturbation theory (DFPT) as implemented in VASP. For these calculations, we used an 8×8×1 k-point mesh on a 3×3 supercell with the PBE-GGA functional. The computed dielectric constants of pristine 1L-SiC were for the in-plane component and for the out-of-plane component, yielding an isotropic average of , which was used in the FNV correction scheme. For the KO correction method, we employed the full dielectric tensor to account for the anisotropic screening effects in the 2D material.
Migration pathways and diffusion barriers () of defect sites were computed using the climbing-image nudged elastic band (CI-NEB) method,49 as presented in Figure 5. This method provides an efficient way to find minimum energy paths between known initial and final states, and accurately determines the transition state energy and configuration. We created seven intermediate images to sample the minimum energy path between the initial and final configurations, which was sufficient to obtain smooth energy profiles and accurately locate transition states. The initial and final states for each migration pathway were fully relaxed before constructing the NEB path. For each vacancy type (Si and C), we considered three distinct migration mechanisms, labeled as Paths A, B, and C, representing different possible routes for vacancy diffusion in the 2D SiC lattice. Path A corresponds to migration to the nearest equivalent site within a single hexagon, Path B involves migration across two hexagons, and Path C represents migration to the nearest neighbor of the same atom type. The spring constant between images was set to 5.0 eV/², which provided adequate coupling between images while allowing sufficient flexibility for the path to converge to the minimum energy route. The climbing image algorithm was employed to precisely locate the transition state by maximizing its energy along the reaction coordinate while minimizing it in all other directions. Force convergence for the NEB calculations was set to 0.05 eV/, which is slightly less stringent than for structural relaxations but sufficient for accurate barrier determination.
For the optical properties, we calculated the frequency-dependent dielectric function within the independent particle approximation using DFPT. This approach, while neglecting excitonic effects, has been shown to provide reliable trends in optical responses for defect systems.11 The imaginary part was obtained by summing over empty states according to:
where and refer to conduction and valence band states respectively, are the cell periodic parts of the wavefunctions at point , are the k-point weights, and is the cell volume. The real part was obtained using the Kramers-Kronig relations:
where denotes the principal value of the integral. From and , we calculated the frequency-dependent absorption coefficient:
where is the speed of light. The optical properties were calculated using DFPT within the PBE framework, which provides a computationally efficient approach for excited state properties while maintaining good accuracy for relative changes in optical response. To ensure accuracy of our optical calculations, we employed a dense 12×12×1 k-point mesh, significantly finer than that used for ground state properties. The calculations included 100 empty bands, chosen after convergence testing showed variations below 0.1% in the computed dielectric function when increasing to 120 bands. We applied a Gaussian broadening of 0.1 eV to the optical spectra, balancing spectral resolution with numerical stability. Local field effects were fully incorporated in our calculations through the inclusion of off-diagonal elements of the dielectric matrix up to 100 eV in the plane-wave basis. The frequency-dependent dielectric response was computed up to 5 eV using a dense frequency grid of 0.02 eV to accurately capture fine spectral features. The absorption coefficients () were computed from the complex dielectric function and are presented as percentages relative to the pristine case to facilitate direct comparison of vacancy effects. The energy range was chosen to focus on the far-infrared region (0-0.5 eV) where vacancy-induced features are most prominent.
3 Results and Discussion
3.1 Defect Formation Energies
Figure 1 presents the analysis of vacancy formation energies in monolayer SiC under carbon-rich conditions, calculated using a supercell with convergence testing performed up to cells. The formation energies are plotted as a function of the Fermi level () for both silicon vacancies (Fig. 1(a)) and carbon vacancies (Fig. 1(b)). These energies were calculated using four distinct exchange-correlation functionals: PBE, SCAN, r2SCAN, and HSE06. The figure displays results obtained using two charge correction schemes: the Freysoldt-Neugebauer-Van de Walle (FNV) method47, shown with dashed lines, and the Kumagai-Oba (KO) method48, represented by solid lines. The range of Fermi levels shown corresponds to the thermodynamic stability region under C-rich conditions.
For silicon vacancies (Fig. 1(a)), the formation energies calculated at the valence band maximum (VBM, ) show a pronounced dependence on the chosen functional. Using the KO correction scheme (solid lines), the energies range from approximately 4.3 eV for PBE (lightest blue) to about 6.0 eV for HSE06 (darkest blue). This significant variation of approximately 2.8 eV underscores the influence of the exchange-correlation functional on the calculated defect energetics. The meta-GGA functionals, SCAN and r2SCAN, yield intermediate results, predicting formation energies at the VBM of around 5.1 eV and 5.5 eV, respectively, with the KO correction. A clear trend is observed where the formation energy systematically increases in the order PBE <SCAN <r2SCAN <HSE06, suggesting that more sophisticated treatments of exchange and correlation stabilize the perfect crystal structure relative to the defected one. As the Fermi level moves towards the conduction band minimum (CBM), the formation energies for silicon vacancies decrease sharply across all functionals, indicating a strong sensitivity to the electronic environment and suggesting that Si vacancies become more energetically favorable under n-type conditions. The charge transition levels, which represent the Fermi energy values where different charge states become the most stable, are found at consistent positions relative to the VBM across the different functionals. Specifically, for Si vacancies (Fig. 1(c)), the (+1/0) transition level occurs at approximately 1.0-1.2 eV above the VBM for HSE06, SCAN, and SCAN, while PBE predicts a notably lower value of 0.43 eV. This discrepancy likely arises from PBE’s underestimation of the band gap and its tendency to overdelocalize defect states, leading to shallower transition levels. The (0/-1) transition level is found at approximately 2.2-2.3 eV for HSE06, 1.37 eV for SCAN, 1.69 eV for SCAN, and 0.84 eV for PBE, as seen in Fig. 1(c). The slopes of the formation energy lines for the charged states ( and ) are exactly and , respectively, consistent with the theoretical formulation (Equation 1) and reflecting the physical process of charge transfer. These transition levels suggest that silicon vacancies can behave as both donors and acceptors within 1L-SiC. Figure 1(c) provides a bar chart summarizing these charge transition levels for Si vacancies across the four functionals.
Carbon vacancies (Fig. 1(b)) consistently exhibit lower formation energies compared to silicon vacancies across all functionals and charge states shown. Using the KO correction (solid lines) at the VBM, the formation energies for C vacancies range from about 4.0 eV (PBE) to about 5.1 eV (HSE06). This lower energy cost compared to Si vacancies (e.g., approximately 5.1 eV for vs 6.0 eV for using HSE06 with KO at the VBM) reflects the relative energetics of removing a C versus a Si atom. The dependence of the formation energy on the functional follows a similar pattern to that observed for Si vacancies, with PBE predicting the lowest energies and HSE06 the highest.The charge transition levels for C vacancies (Fig. 1(d)) show the (+1/0) transition occurring near 0.8-0.9 eV above the VBM, and the (0/-1) transitions occurring around 1.1 eV for PBE, SCAN, and r2SCAN, and around 1.6-1.7 eV for HSE06. However, an interesting difference is that the formation energy of carbon vacancies decreases less steeply with an increasing Fermi level compared to silicon vacancies. For example, using HSE06 (KO), the formation energy for C vacancies decreases by about 1.3 eV over the first eV increase in Fermi level, compared to a steeper 1.8 eV decrease for Si vacancies over the same range. This suggests a reduced sensitivity to the electronic environment for C vacancies and implies they might be less effective as compensating defects in n-type conditions compared to Si vacancies. Figure 1(d) presents a bar chart illustrating these charge transition levels for C vacancies for each functional.
Comparing the charge correction schemes, the Kumagai-Oba (KO) method48 (solid lines) consistently yields slightly higher formation energies (by 0.1-0.2 eV) than the Freysoldt-Neugebauer-Van de Walle (FNV) method47 (dashed lines) across most functionals and Fermi levels shown in the figure. This discrepancy stems from the KO method’s more thorough treatment of the anisotropic dielectric screening inherent in 2D materials, utilizing the explicitly calculated in-plane () and out-of-plane () dielectric constants. The difference between the correction schemes underscores the importance of employing appropriate charge correction methods for accurate defect calculations in 2D systems. The thermodynamic stability window, indicating the range of Fermi levels where these vacancies can exist in thermal equilibrium, spans from the VBM up to approximately 1.6-1.7 eV (the(0/-1) transition level for HSE06) for C vacancies, and up to approximately 2.2-2.3 eV (the (0/-1) transition level for HSE06) for Si vacancies under the assumed carbon-rich conditions, when considering these transition levels as the upper limits within the displayed energy range of the figure. This range is constrained by the formation of competing defects and the material’s intrinsic band gap, although the absolute values of the formation energies are dependent on the specific chemical potential references chosen for silicon and carbon.
3.2 Electronic Properties
[…]